2y - x =3
x = 3y – 5
Solve by substitution method.

Answer by KnightOwlTutor(293)

2y - x =3
x = 3y – 5
Replace x in the first equation with 3y-5
2y-(3y-5)=3
2y-3y+5=3
-y+5=3
Subtract 5 from both sides
-y=-2
Divide both sides by -1
y=2

x=3(2)-5
x=1
Let's test our answer 2(2)-1=3
4-1=3